# AI-Slop to AI-Polish? Aligning Language Models through Edit-Based Writing Rewards and Test-time Computation

### v1

@@ Abstract
AI-generated text is proliferating across domains, from creative writing and journalism to marketing content and scientific articles. Models can follow user-provided instructions to generate coherent and grammatically correct outputs but in this work, we study a more fundamental question: how do we evaluate and improve the writing quality of AI-generated text? Writing quality assessment has received less attention from the community, in part because it is fundamentally subjective and requires expertise. We first introduce the Writing Quality Benchmark (WQ) by consolidating five writing-preference datasets into 4,729 writing quality judgments. Our experiments show that competitive baselines, including state-of-the-art LLMs that excel at reasoning tasks, barely outperform random baselines on WQ. We then train specialized Writing Quality Reward Models (WQRM) of various sizes for writing quality assessment that demonstrate strong generalization on four out-of-distribution test sets and 74% accuracy on the WQ benchmark. To further show WQRM’s practical benefits during inference, we leverage additional test-time compute to generate and rank multiple candidate revisions, allowing us to select higher-quality outputs from an initial draft. Human evaluation with 9 experienced writers confirm that WQRM-based selection produces writing samples preferred by experts 66% overall, and 72.2% when the reward gap is larger than 1 point. We release our datasets and models to encourage community engagement with writing quality assessment and development of AI writing systems better aligned with human preferences.
@@ 1 Introduction
Writing is one of the most important pillars of education, enabling learners to critically engage with the topics they study. In The Rise of Writing Brandt (2014) argues that the “information economy’s insatiable demand for symbol manipulation—‘knowledge work’—has forced many workers to reorient their labor around the production of prose” (Laquintano & Vee, 2024). Generative AI tools have further blurred these boundaries, especially around how labor and writing practices are evolving across both academic (Kobak et al., 2024; Lee et al., 2025) and professional contexts (Liang et al., 2025). Often awkward and jarring to read, low-effort text generated by AI is now flooding web browsers and social-media platforms much like spam in old inboxes (Herrman, 2024a; Knibbs, 2024c; d; b; a). This neologistic term of revulsion is often referred to as “A.I. slop” (Herrman, 2024b). Extensive social experimentation with ChatGPT has invited criticism on social media and in the popular news platforms that its writing has a disembodied “robovoice”. This has led to humanization methods (Wang et al., 2024) and even start-ups such as StealthGPT  [^1] https://www.stealthgpt.ai/ or HumanizeAI  [^2] https://www.humanizeai.pro/, which explicitly attempt to make AI-generated text more humanlike.
Despite LLMs showing impressive performance in math and coding, their ability to write high-quality text has been rather pedestrian. Recent work from Chakrabarty et al. (2024b) shows how text generated from widely used LLMs are often rife with clichés, purple prose, poor sentence structure, and unnecessary exposition. This stems from several challenges. Unlike math or coding, writing lacks verifiable rewards. While it would be possible to train a model to write better text by having humans label examples of “good” and “bad” writing, it is challenging due to the required expertise. Self-evaluation using LLMs has proven useful in reward modeling and constitutional AI (Bai et al., 2022), but relying on uncalibrated humans or LLMs for feedback (Lee et al., 2023; Gao et al., 2024) on subjective tasks like writing can lead to reward hacking (Pan et al., 2024) and alignment issues. Recent work from Panickssery et al. (2024) shows the self-aggrandizing nature of LLMs, as evidenced in Table 3 in the Appendix B, where they prefer their own writing over Nobel Prize winners’ work.
@@
The surge in AI writing assistance demands urgent alignment of AI-generated text with human preferences. Recent work from Gooding et al. (2025) show how LLMs struggle to select high-quality writing actions as judged by human experts, often treating suboptimal and optimal interventions as equally acceptable. They highlight the need for models to better assess the quality and impact of suggested actions, both during generation and across multi-step refinement. Binary preference feedback between paired examples is the most common alignment method for LLMs (Christiano et al., 2017), but it has a significant drawback. The paired outputs may differ in several ways and could be equally worse in terms of quality (Casper et al., 2023; Lambert & Calandra, 2023). [^3] Forcing annotators to choose between two undesirable outputs doesn’t improve alignment. In the current design of RLHF, annotators are not allowed to pick neither Recent work from Chakrabarty et al. (2024b) shows how identifying and editing problematic response segments effectively improves AI alignment. This also reflects the Reviewing phase in the cognitive process model of writing (Hayes et al., 1987), where humans evaluate and revise text. They release LAMP (Language model Authored, Manually Polished), a corpus of 1282 $<AI-generated,Expert-Edited>$ pairs with implicit preference ($edited>original\_draft$) to improve AI writing (see Table 4 in Appendix C). Additionally, each paragraph pair includes normalized scores (1-10) reflecting writing quality before and after editing.
@@ 3 Writing Quality Reward Models
We rely on the LAMP (Language model Authored, Manually Polished) corpus from Chakrabarty et al. (2024b) to train reward models. As illustrated in Figure 2, each sample in LAMP consists of a writing instruction and two paragraphs that match this instruction. The paragraphs in LAMP range from 150 to 400 words, and span across fiction and non-fiction. Table 4 in Appendix C shows a sample from LAMP, highlighting the edits implemented by an expert to improve writing quality. We use three methods to transform LAMP samples into training and validation data points for our models: pairwise (P), scalar (R), and combined (PR). With the P method, each data point presents two paragraphs as input (1 and 2) and requires a binary classification output indicating which paragraph has higher writing quality (i.e., the output is 1 or 2). Each LAMP sample is duplicated into two P data points by considering both paragraph orders (AI-generated, Expert-Edited $\to$ 2) and (Expert-Edited, AI-generated $\to$ 1). With the R method, each data point takes a single paragraph as input and outputs a regression value predicting the quality score of that paragraph. Since each LAMP sample contains two paragraphs (before and after edit), it generates two R data points. The PR method combines both approaches, yielding four data points per LAMP sample (two from P and two from R). There are a total of 1,282 samples in LAMP, and we follow the author’s split divisions of 1,000 training, 67 validation, and 215 test samples. Applying the data transformation described above, the P, R, and PR variants of the training data we obtain consist of 2,000, 2,000, and 4,000 training data points, respectively. For our experiments, we trained both generative LLMs (Llama3.1 (Dubey et al., 2024)) and encoder-only models (ModernBert (Warner et al., 2024)).
@@
A key advantage of generative methods is their potential to produce natural-language rationales alongside their predictions, which can be leveraged for the interpretability of the model responses. Prior work from Wiegreffe et al. (2020) has demonstrated label-rationale association as an important property for faithfulness of model response. Additionally researchers (Ludan et al., 2023; Hase & Bansal, 2021) have argued for incorporating explanations in both model input or output as way to make it robust against spurious cues. Since the original LAMP corpus did not collect rationales from experts during annotation, we introduce a procedure to augment LAMP with LLM-generated silver rationales. Specifically, we first collected five examples that either contrast one paragraph’s strength over the other (P-style) or holistically critique or praise the paragraph (R-style) from two professional writers. Writers were specifically instructed to cite excerpts from the paragraphs to justify their rationales. We then use these expert-written rationales as demonstration examples in an in-context instruction and generate rationales using the Claude3.5 Sonnet model  [^5] Considered a top-performing model for writing tasks at the time of experiments.. The Appendix’s Table 6 in Appendix E shows examples of rationale generated by Claude3.5 Sonnet.
@@ 4 The Writing Quality Benchmark
To assess writing quality prior work has argued for evaluation by professionals (ones with writing experience). Nevertheless, some writing quality preference datasets are based on crowd-sourced judgments. We include four datasets based on expert judgments and one dataset based on crowd-sourced annotation (LM Arena) to represent both perspectives in the benchmark. Finally, we selected two datasets with long responses (Art or Artifice, LM Arena) and three with shorter responses ranging from 200-400 words, ensuring that models that perform well on the benchmark are capable of judging writing quality irrespective of length. Appendix E.1 details the procedure we followed to extract and standardize each dataset.
@@ 4.1 Experimental Results on WQ
Experimental results are summarized in Table 2. First, we find that all the LLMs used in zero-shot settings perform below or a few percentage points above a random baseline of 50%. The performance is particularly low on portions of WQ that involve AI-human preference pairs. This confirms prior findings that LLMs used in LLM-as-a-judge settings tend to prefer AI-generation over human-writing (Panickssery et al., 2024). The O1 and R1 reasoning models do not significantly outperform their non-reasoning counterparts, indicating that out-of-the box COT-style reasoning, useful for math or coding tasks doesn’t improve writing quality assessement. Finally, adding five few-shot examples to GPT-4o does help improve performance from 40.9 to 54.3, however further experiments with additional in-context examples did not lead to further gains, confirming that few-shot examples in the instruction are not sufficient to achieve strong performance on WQ.
@@ 5.1 Generating edits via Supervised Finetuning
Prior work from Chakrabarty et al. (2024b) shows experimentally that LLMs’ text idiosyncrasies (cliches, redundancy, lack of subtext, etc.) can be mitigated through self-editing in an in-context setup. Borrowing motivation from them we teach LLMs how to improve their response via edits. Figure 3 illustrates the three components of the editing pipeline. Given a first draft response to an instruction from any given LLM, the first step consists of identifying and listing idiosyncrasies: spans in the first draft that can be rephrased to improve overall writing quality. For each identified idiosyncrasy, a second stage consists in rewriting the idiosyncrasy. This is framed as an executable edit (Laban et al., 2023), where each edit consists of replacing an original string in a draft with an improved version. The third step simply executes all edits (by applying a series of string replace operations) to obtain the final edited draft. While Chakrabarty et al. (2024b) implemented this through prompt-chaining (Wu et al., 2022) with few-shot examples, we improved efficiency by supervised fine-tuning of GPT-4o and Llama3.1 70B based on the entire LAMP training set. The training input consists of the first draft alongside the entire edit interaction trace (detect, rewrite, execute) in a step-by-step chain of thought prompt, and the output is the edited paragraph. See Appendix G for an example COT prompt.
@@ 6 How well calibrated are our reward models ?
We generated 100 draft responses (50 GPT4-o, 50 Llama3.1 70B) based on 90 writing instructions spanning 3 domains: literary fiction, non-fiction, and product marketing. For literary fiction and non-fiction we create the instructions through instruction back-translation (Li et al., 2023) conditioned on expert-written paragraphs in Anonymous (2025) and news articles in the data from Russell et al. (2025). Marketing writing instructions were based on products recommended in WireCutter articles [^7] https://www.nytimes.com/wirecutter/ across the Home, Kitchen and Tech sections. The right portion of Figure  7 summarizes the process we follow to leverage test-time compute. Specifically, we obtain a first draft from a LLM (GPT4o or Llama3.1 70B) followed by drawing $N=20$ candidate edited responses from the respective SFT model (Section 5.1)  [^8] If first draft is from GPT4o we use GPT4o SFT model, and score each candidate with the WQRM model. We filter out any candidate that scores lower than the first drafts, and then form response triplets by selecting the first draft, a randomly-selected edited response (random edit), and the Best-of-N candidate response according to WQRM (Best Edit) (See example triplet in Table 9). We recruited 9 professional writers through mailing lists from top MFA programs in the US. They were asked to rank three responses based on its overall quality (See Figure 7 for interface). Each response triplet were annotated by three experts, which we aggregated into a majority rank. Participants completed annotation in batches of 10 triplets at a time, and were paid \$100 per batch.
@@ 6.1 Study Findings
Figure 4 summarizes findings from the expert annotation. In Figure 4(a), we plot the distribution of rankings across all triplets. Best Edit candidates were most preferred overall with an average rank of 1.58, followed by random edit (2.09) and first draft (2.26). The breakdown of rankings across domains (fiction, non-fiction, marketing) or LLM (GPT-4o vs. Llama 3.1) is presented in Appendix H. In short, Best Edit achieves the top rank in all conditions, confirming the generalization of WQRM scores across conditions.
@@ Appendix A Appendix
Appendix B Motivation Example
Which is better quality writing?
Appendix C Example LAMP Sample
@@
Appendix D Generative WRQM Prompt Formats
@@
Appendix E Example Rationales For Generative WQRM Training
@@ E.1 Datasets
Appendix F Example Human Mimic Samples
@@
Appendix G Example COT Editing Prompt
@@
Appendix H Expert Annotation Result Breakdown
@@ Appendix I Comparison
Appendix J Expert Annotation Interface

### v2

@@ Abstract
AI-generated text is proliferating across domains, from creative writing and journalism to marketing content and scientific articles. Models can follow user-provided instructions to generate coherent and grammatically correct outputs but in this work, we study a more fundamental question: how do we evaluate and improve the writing quality of AI-generated text? Writing quality assessment has received less attention from the community, in part because it is fundamentally subjective and requires expertise. We first introduce the Writing Quality Benchmark (WQ) by consolidating five writing-preference datasets into 4,729 writing quality judgments. Our experiments show that most of the competitive baselines, including state-of-the-art LLMs that excel at reasoning tasks, barely outperform random baselines on WQ. We then train specialized Writing Quality Reward Models (WQRM) of various sizes for writing quality assessment that demonstrate strong generalization on four out-of-distribution test sets and 74% accuracy on the WQ benchmark. To further show WQRM’s practical benefits during inference, we leverage additional test-time compute to generate and rank multiple candidate revisions, allowing us to select higher-quality outputs from an initial draft. Human evaluation with 9 experienced writers confirm that WQRM-based selection produces writing samples preferred by experts 66% overall, and 72.2% when the reward gap is larger than 1 point. We release our datasets and models to encourage community engagement with writing quality assessment and development of AI writing systems better aligned with human preferences.
@@ 1 Introduction
Writing is one of the most important pillars of education, enabling learners to critically engage with the topics they study. In The Rise of Writing Brandt (2014) argues that the “information economy’s insatiable demand for symbol manipulation—‘knowledge work’—has forced many workers to reorient their labor around the production of prose” (Laquintano & Vee, 2024). Generative AI tools have further blurred these boundaries, especially around how labor and writing practices are evolving across both academic (Kobak et al., 2024; Lee et al., 2025) and professional contexts (Liang et al., 2025). Often awkward and jarring to read, low-effort text generated by AI is now flooding web browsers and social-media platforms much like spam in old inboxes (Herrman, 2024a; Knibbs, 2024c; d; b; a). This neologistic term of revulsion is often referred to as “A.I. slop” (Herrman, 2024b). Extensive social experimentation with ChatGPT has invited criticism on social media and in the popular news platforms that its writing has a disembodied “robovoice”. This has led to humanization methods (Wang et al., 2024) and even start-ups such as StealthGPT or HumanizeAI, which explicitly attempt to make AI-generated text more humanlike.
Despite LLMs showing impressive performance in math and coding, their ability to write high-quality text has been rather pedestrian. Recent work from Chakrabarty et al. (2024b) shows how text generated from widely used LLMs are often rife with clichés, purple prose, poor sentence structure, and unnecessary exposition. This stems from several challenges. Unlike math or coding, writing lacks verifiable rewards. While it would be possible to train a model to write better text by having humans label examples of “good” and “bad” writing, it is challenging due to the required expertise. Self-evaluation using LLMs has proven useful in reward modeling and constitutional AI (Bai et al., 2022), but relying on uncalibrated humans or LLMs for feedback (Lee et al., 2023; Gao et al., 2024) on subjective tasks like writing can lead to reward hacking (Pan et al., 2024) and alignment issues. Recent work from Panickssery et al. (2024) shows the self-aggrandizing nature of LLMs, as evidenced in Table 1 where they prefer their own writing over Nobel Prize winners’ work.
@@
The surge in AI writing assistance demands urgent alignment of AI-generated text with human preferences. Recent work from Gooding et al. (2025) show how LLMs struggle to select high-quality writing actions as judged by human experts, often treating suboptimal and optimal interventions as equally acceptable. They highlight the need for models to better assess the quality and impact of suggested actions, both during generation and across multi-step refinement. Binary preference feedback between paired examples is the most common alignment method for LLMs (Christiano et al., 2017), but it has a significant drawback. The paired outputs may differ in several ways and could be equally worse in terms of quality (Casper et al., 2023; Lambert & Calandra, 2023). [^1] Forcing annotators to choose between two undesirable outputs doesn’t improve alignment. In the current design of RLHF, annotators are not allowed to pick neither Recent work from Chakrabarty et al. (2024b) shows how identifying and editing problematic response segments effectively improves AI alignment. This also reflects the Reviewing phase in the cognitive process model of writing (Hayes et al., 1987), where humans evaluate and revise text. They release LAMP (Language model Authored, Manually Polished), a corpus of 1282 $<AI-generated,Expert-Edited>$ pairs with implicit preference ($edited>original\_draft$) to improve AI writing (see Table 4 in Appendix B). Additionally, each paragraph pair includes normalized scores (1-10) reflecting writing quality before and after editing.
Which is better quality writing?
@@ 3 Writing Quality Reward Models
We rely on the LAMP (Language model Authored, Manually Polished) corpus from Chakrabarty et al. (2024b) to train reward models. As illustrated in Figure 2, each sample in LAMP consists of a writing instruction and two paragraphs that match this instruction. The paragraphs in LAMP range from 150 to 400 words, and span across fiction and non-fiction. Table 4 in Appendix B shows a sample from LAMP, highlighting the edits implemented by an expert to improve writing quality. We use three methods to transform LAMP samples into training and validation data points for our models: pairwise (P), scalar (R), and combined (PR). With the P method, each data point presents two paragraphs as input (1 and 2) and requires a binary classification output indicating which paragraph has higher writing quality (i.e., the output is 1 or 2). Each LAMP sample is duplicated into two P data points by considering both paragraph orders (AI-generated, Expert-Edited $\to$ 2) and (Expert-Edited, AI-generated $\to$ 1). With the R method, each data point takes a single paragraph as input and outputs a regression value predicting the quality score of that paragraph. Since each LAMP sample contains two paragraphs (before and after edit), it generates two R data points. The PR method combines both approaches, yielding four data points per LAMP sample (two from P and two from R). There are a total of 1,282 samples in LAMP, and we follow the author’s split divisions of 1,000 training, 67 validation, and 215 test samples. Applying the data transformation described above, the P, R, and PR variants of the training data we obtain consist of 2,000, 2,000, and 4,000 training data points, respectively. For our experiments, we trained both generative LLMs (Llama3.1 (Dubey et al., 2024)) and encoder-only models (ModernBert (Warner et al., 2024)).
@@
A key advantage of generative methods is their potential to produce natural-language rationales alongside their predictions, which can be leveraged for the interpretability of the model responses. Prior work from Wiegreffe et al. (2020) has demonstrated label-rationale association as an important property for faithfulness of model response. Additionally researchers (Ludan et al., 2023; Hase & Bansal, 2021) have argued for incorporating explanations in both model input or output as way to make it robust against spurious cues. Since the original LAMP corpus did not collect rationales from experts during annotation, we introduce a procedure to augment LAMP with LLM-generated silver rationales. Specifically, we first collected five examples that either contrast one paragraph’s strength over the other (P-style) or holistically critique or praise the paragraph (R-style) from two professional writers. Writers were specifically instructed to cite excerpts from the paragraphs to justify their rationales. We then use these expert-written rationales as demonstration examples in an in-context instruction and generate rationales using the Claude3.5 Sonnet model  [^3] Considered a top-performing model for writing tasks at the time of experiments.. The Appendix’s Table 6 in Appendix D shows examples of rationale generated by Claude3.5 Sonnet.
@@ 4 The Writing Quality Benchmark
To assess writing quality prior work has argued for evaluation by professionals (ones with writing experience). Nevertheless, some writing quality preference datasets are based on crowd-sourced judgments. We include four datasets based on expert judgments and one dataset based on crowd-sourced annotation (LM Arena) to represent both perspectives in the benchmark. Finally, we selected two datasets with long responses (Art or Artifice, LM Arena) and three with shorter responses ranging from 200-400 words, ensuring that models that perform well on the benchmark are capable of judging writing quality irrespective of length. Appendix D.1 details the procedure we followed to extract and standardize each dataset.
@@ 4.1 Experimental Results on WQ
Experimental results are summarized in Table 3. First, we find that all the LLMs used in zero-shot settings perform below or a few percentage points above a random baseline of 50%. The performance is particularly low on portions of WQ that involve AI-human preference pairs. This confirms prior findings that LLMs used in LLM-as-a-judge settings tend to prefer AI-generation over human-writing (Panickssery et al., 2024). The O1 and R1 reasoning models do not significantly outperform their non-reasoning counterparts, indicating that out-of-the box COT-style reasoning, useful for math or coding tasks doesn’t improve writing quality assessement. O3 shows improvement on Synthetic Mirror and Art or Artifice showing some promise.Finally, adding five few-shot examples to GPT-4o does help improve performance from 40.9 to 54.3, however further experiments with additional in-context examples did not lead to further gains, confirming that few-shot examples in the instruction are not sufficient to achieve strong performance on WQ.
@@ 5.1 Generating edits via Supervised Finetuning
Prior work from Chakrabarty et al. (2024b) shows experimentally that LLMs’ text idiosyncrasies (cliches, redundancy, lack of subtext, etc.) can be mitigated through self-editing in an in-context setup. Borrowing motivation from them we teach LLMs how to improve their response via edits. Figure 3 illustrates the three components of the editing pipeline. Given a first draft response to an instruction from any given LLM, the first step consists of identifying and listing idiosyncrasies: spans in the first draft that can be rephrased to improve overall writing quality. For each identified idiosyncrasy, a second stage consists in rewriting the idiosyncrasy. This is framed as an executable edit (Laban et al., 2023), where each edit consists of replacing an original string in a draft with an improved version. The third step simply executes all edits (by applying a series of string replace operations) to obtain the final edited draft. While Chakrabarty et al. (2024b) implemented this through prompt-chaining (Wu et al., 2022) with few-shot examples, we improved efficiency by supervised fine-tuning of GPT-4o and Llama3.1 70B based on the entire LAMP training set. The training input consists of the first draft alongside the entire edit interaction trace (detect, rewrite, execute) in a step-by-step chain of thought prompt, and the output is the edited paragraph. See Appendix F for an example COT prompt.
@@ 6 How well calibrated are our reward models ?
We generated 100 draft responses (50 GPT4-o, 50 Llama3.1 70B) based on 90 writing instructions spanning 3 domains: literary fiction, non-fiction, and product marketing. For literary fiction and non-fiction we create the instructions through instruction back-translation (Li et al., 2023) conditioned on expert-written paragraphs in Anonymous (2025) and news articles in the data from Russell et al. (2025). Marketing writing instructions were based on products recommended in WireCutter articles [^5] https://www.nytimes.com/wirecutter/ across the Home, Kitchen and Tech sections. The right portion of Figure  1 summarizes the process we follow to leverage test-time compute. Specifically, we obtain a first draft from a LLM (GPT4o or Llama3.1 70B) followed by drawing $N=20$ candidate edited responses from the respective SFT model (Section 5.1)  [^6] If first draft is from GPT4o we use GPT4o SFT model, and score each candidate with the WQRM model. We filter out any candidate that scores lower than the first drafts, and then form response triplets by selecting the first draft, a randomly-selected edited response (random edit), and the Best-of-N candidate response according to WQRM (Best Edit) (See example triplet in Table 9). We recruited 9 professional writers through mailing lists from top MFA programs in the US. They were asked to rank three responses based on its overall quality (See Figure 7 for interface). Each response triplet were annotated by three experts, which we aggregated into a majority rank. Participants completed annotation in batches of 10 triplets at a time, and were paid \$100 per batch.
@@ 6.1 Study Findings
Figure 4 summarizes findings from the expert annotation. In Figure 4(a), we plot the distribution of rankings across all triplets. Best Edit candidates were most preferred overall with an average rank of 1.58, followed by random edit (2.09) and first draft (2.26). The breakdown of rankings across domains (fiction, non-fiction, marketing) or LLM (GPT-4o vs. Llama 3.1) is presented in Appendix G. In short, Best Edit achieves the top rank in all conditions, confirming the generalization of WQRM scores across conditions.
@@ Appendix A Appendix
Appendix B Example LAMP Sample
@@
Appendix C Generative WRQM Prompt Formats
@@
Appendix D Example Rationales For Generative WQRM Training
@@ D.1 Datasets
Appendix E Example Human Mimic Samples
@@
Appendix F Example COT Editing Prompt
@@
Appendix G Expert Annotation Result Breakdown
@@ Appendix H Comparison
Appendix I Expert Annotation Interface